# Embedding Space Allocation with Angle-Norm Joint Classifiers for Few-Shot Class-Incremental Learning

### v1

@@ 4.1.2 Feature alignment
We design a novel Cosine Center (CC) loss to guide sample embeddings closer to their corresponding class centers and away from other class centers in the angular space, achieving the effect of feature alignment. We define the CC loss $\mathcal{L}_{\text{cc}}$ as $\alpha\mathcal{L}_{1}+\beta\mathcal{L}_{2},$ where $\mathcal{L}_{1}$ pulls embeddings closer to their corresponding centers, $\mathcal{L}_{2}$ pushes embeddings away from other class centers, and $\alpha$ and $\beta$ are weighting coefficients. $\mathcal{L}_{1}$ is defined by:
@@ 4.1.3 Center updates
The pre-assigned and fixed centers may not semantically match the classes. In other words, centers that are semantically close may be far apart in embedding space. To ensure the semantic coherence of the centers, we combine CE loss and CC loss to jointly guide the model and adjust the centers during the learning process. The total loss is defined as:
@@
where $\hat{\mathbf{e}}_{i}=\frac{\mathbf{e}_{i}}{\|\mathbf{e}_{i}\|}$. Continuously adding the mean of normalized embeddings to the center will gradually move the center closer to the sample embeddings which learned by CE loss. Setting the centers as learnable parameters and updating them through gradient descent is equivalent to using Eq. 6, as shown in A.
@@ 4.2.1 Log-Norm distribution of embeddings
If the norms of samples from different categories have distinct distributions, it indicates that norms contain label information to a certain extent. The distributions of $\ln\|\mathbf{e}\|$ from different sessions and categories, as extracted by a CNN trained in the FSCIL paradigm, are shown in Fig. 4. In Fig. 4(a), it can be observed that the mean and variance of embedding log-norms during the base sessions are significantly higher than those during the incremental sessions. This is may due to the fact that in FSCIL, there are only a few samples in the incremental sessions, resulting in the model’s limited ability to fit new samples. Additionally, as shown in Fig. 4(b), there are differences in the norms of different categories within the same session, especially in session 0, which may be caused by the model’s varying degrees of fit to different categories. Although there is significant overlap in the norms across different categories and sessions, preventing direct prediction based solely on norms, norms can still serve as a part of the basis for classification.
Figure 4: Histogram of the distribution of the log-norms of sample embedding on miniImageNet dataset. (a) 3 different sessions. (b) 3 randomly selected classes in the base session and 3 in incremental sessions. The curve represents the normal distribution formed by the sample mean and variance of the log-norms.
@@ 4.2.2 Dual-Granularity parametric estimation
As shown in Fig. 4, we observe significant differences between the incremental and base sessions, but the distributions across different incremental sessions are approximately similar. Based on this phenomenon, we choose to estimate a distribution for the norms of each class. Based on the results of the observation, we assume that the log-norms of the embeddings for samples belonging to the same category follow a normal distribution, represented as $\ln{\|\mathbf{e}_{i}\|}\sim\mathcal{N}(\mu_{j},\sigma^{2}_{j})$, where $y_{i}=j$, $j$ is the category label, $\mu_{j}$ denotes the mean, and $\sigma^{2}_{j}$ denotes the variance.
During the base session, with sufficient training samples, we approximate the population mean and variance with sample mean and variance, respectively. As shown in Fig. 4(b), the differences in log-norms between classes are significant, making it necessary to perform class-level modeling of the log-norms. During the incremental sessions, however, each class contains only a few samples, which may result in significant deviations between the sample mean and variance and the true population values. Despite this, as shown in Fig. 4(b), the differences in log norm between classes during these sessions are actually quite small. We assume that the log-norms of all incremental categories follow the same normal distribution, denoted as $\ln{\|\mathbf{e}_{i}\|}\sim\mathcal{N}(\mu_{\text{incre}},\sigma^{2}_{\text{incre}})$, where $y_{i}\in(\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t})$, $t\geq 1$, and $\mu_{\text{incre}},\sigma^{2}_{\text{incre}}$ are the shared mean and variance, respectively. In summary, we model the distributions of the base and incremental rounds with different levels of granularity.
@@ 4.2.3 Angle-Norm joint logits inference
The pipeline of Angle-Norm joint (ANJ) inference is illustrated in Fig. 3. ANJ consists of 2-Stage NCM (2SNCM) and Norm Distribution (ND).
@@
where $\hat{\mathbf{e}}_{i}=\frac{\mathbf{e}_{i}}{\|\mathbf{e}_{i}\|}$. During the base session, we normalize embeddings to treat each sample equally, because otherwise, samples with larger norms would have greater weight in the calculation of $\boldsymbol{w}_{j}$. As illustrated in Fig. 4(a), there is considerable variance in the norms of samples during the base session. In the incremental sessions, due to the scarcity of samples, normalizing embeddings could lead to significant loss of information. Therefore, we adopt a staged approach to better accommodate FSCIL.
@@
where $X$ represents a random variable that follows the distribution of norms for class $j$. We can obtain the probability by cumulative distribution function of normal distribution.
Angle-Norm Joint Inference. Since norms contain less information than angles, we first compress the norm logits, ultimately obtaining the joint logits. The final angle-norm joint logits $\mathbf{z}_{j}$ is calculated by $\mathbf{z}_{j}=\mathbf{z}_{j}^{1}\left(\mathbf{z}_{j}^{2}\right)^{C}$, where $C$ is the compression coefficient. In our experiments, $C$ is set to 0.006. After compression, the differences between the norm logits of different categories are reduced, thereby amplifying the role of the angle logits. The final predicted category is determined by $\mathbf{z}_{j}$.
@@ 5.1 Experimental Setup
Datasets. Following the previous works setting [8, 11, 9, 16], we perform experiments on three datasets: CIFAR100 [48], CUB200 [52], and miniImageNet [53]. CIFAR100 consists of 100 different classes, with each class containing 600 images. CUB200 contains images of 200 bird species. miniImageNet is a subset of ImageNet [1] and consists of 100 selected classes. In CIFAR100 and miniImageNet, the base session consists of 60 classes, while the incremental sessions comprise 40 classes. During the incremental sessions, the data is organized in a 5-way 5-shot format. As for CUB200, the base session includes 100 classes, and the incremental sessions involve 100 classes as well. In the incremental sessions of CUB200, the data is organized in a 10-way 5-shot format.
Implementation Details. We conduct experiments by separately employing SAAN, and embed it into SOTA methods, $i.e.$ FACT [11], SAVC [16]. All the three methods adopt the feature freezing paradigm and use NCM classifier. For fairness, all these methods adopt ResNet18 [2] for miniImageNet and CUB200, and ResNet20 [2] for CIFAR100. In SAAN, the initial learning rate is set to 0.1 with cosine annealing on CIFAR100 and miniImageNet, and milestone decay on CUB200. The batch size is 256, and training is conducted for 600 epochs. The training parameters in the embedding of FACT and SAVC are set to be the same as those in their works. The selection of hyper-parameters used in SAAN is discussed in Sec. 5.4.
Table 1: Performance comparison on CUB200 dataset. *: FSCIL results implemented by [8]. $\ddagger$:reproduced under FSCIL setting with the source code. Drop: The drop in Acc. of the last session relative to session 0.
@@
Table 2: Performance comparison on miniImageNet dataset. *: FSCIL results implemented by [8]. $\ddagger$:reproduced under FSCIL setting with the source code. Drop: The drop in Acc. of the last session relative to session 0.
@@
Table 3: Performance comparison on CIFAR100 dataset. *: FSCIL results implemented by [8]. $\ddagger$:reproduced under FSCIL setting with the source code. Drop: The drop in Acc. of the last session relative to session 0.
@@ 5.2 Performance on Benchmarks
Compared to Decoupled-Cosine (frozen paradigm baseline), our method exhibits an average improvement over 7.44% across the three datasets in the final session, highlighting the importance of prior spatial allocation. It can be observed that SAAN surpasses FACT on all three datasets, indicating that our method achieves performance of SOTA models. FACT occupies the sample space with virtual classes, but new classes may not appear within the space occupied by these virtual classes. SAAN further aligns the virtual classes semantically with the new classes, addressing the limitations of FACT, which allows SSAN to achieve better performance.
SAAN does not outperform SAVC, primarily because SAVC not only uses the virtual class method but also incorporates supervised contrastive learning. Based on the self-supervised MoCo [56] framework, it significantly increases both data volume and training capacity with additional supervisory signals, yielding highly effective results. However, the SAAN framework does not conflict with SAVC.
@@ 5.3 Ablation Studies
The experiments reveal that $\mathcal{L}_{1}$ has the most significant impact, with a 3.41% improvement when combined with the baseline, highlighting the importance of embedding space allocation. $\mathcal{L}_{2}$ aims to increase inter-class distances, resulting in a 4.01% improvement. 2S achieves a modest improvement of 4.19%, indicating a limited enhancement. This indicates that 2S can achieve more reasonable representative points simply by normalizing and averaging the samples in the base round, ultimately leading to a certain performance improvement. To further verify the effectiveness of 2S, we present multiple repeated experiments on three datasets in C. ND achieves an improvement of 5.18%, demonstrating the effectiveness of introducing norm information to classifier. The ANJ, composed of 2S and ND, ultimately achieves a performance improvement of 5.44%. The components in ANJ do not participate in model training but construct a more accurate classifier based on the feature extractor, with minimal computational cost. In summary, the spatial allocation strategy of CCSA is highly beneficial for FSCIL, while ANJ aids classification by making more comprehensive and effective use of the features.
@@
Figure 5: The t-SNE visualization on CIFAR100 dataset of the embeddings. Class 0-5 represent the base classes while class 6-8 represent the novel classes. The squares represent the centers set by SAAN.

### v2

@@ 4.1.2 Feature alignment
We design a novel cosine center loss to guide sample embeddings closer to their corresponding class centers and away from other class centers in the angular space, achieving the effect of feature alignment. We define the cosine center loss $\mathcal{L}_{\text{cc}}$ as $\alpha\mathcal{L}_{1}+\beta\mathcal{L}_{2},$ where $\mathcal{L}_{1}$ pulls embeddings closer to their corresponding centers, $\mathcal{L}_{2}$ pushes embeddings away from other class centers, and $\alpha$ and $\beta$ are weighting coefficients. $\mathcal{L}_{1}$ is defined by:
@@ 4.1.3 Center updates
The pre-assigned and fixed centers may not semantically match the classes. In other words, semantically close centers may be far apart in embedding space. To ensure the semantic coherence of the centers, we combine CE loss and cosine center loss to jointly guide the model and adjust the centers during the learning process. The total loss is defined as:
@@
where $\hat{\mathbf{e}}_{i}=\frac{\mathbf{e}_{i}}{\|\mathbf{e}_{i}\|}$. Continuously adding the mean of normalized embeddings to the center will gradually move the center closer to the sample embeddings which is learned by cross-entropy loss. Setting the centers as learnable parameters and updating them through gradient descent is equivalent to using Eq. 6, as shown in A.
@@ 4.2.1 Log-Norm distribution of embeddings
NCM is based on cosine similarity, so it is an algorithm that relies on the angular relationship between vectors while ignoring the vector norm (the distance between the vector’s tip and the origin). We attempt to incorporate this vector length information into the classification criteria. If the norms of samples from different categories have distinct distributions, it indicates that norms contain label information to a certain extent. The distributions of $\ln\|\mathbf{e}\|$ from different sessions and categories, as extracted by a CNN trained in the FSCIL paradigm, are shown in Fig. 4. In Fig. 4(a) and 4(c), it can be observed that the mean and variance of embedding log norms during the base sessions are significantly higher than those during the incremental sessions. Based on the characteristic demonstrated by the embedding log norm, we attempt to leverage this feature to assist us in determining whether this class belongs to the base session or the incremental sessions during inference. Additionally, as shown in Fig. 4(b) and 4(d), in the base session, the log-norm distributions of different classes vary significantly, whereas in the incremental sessions, this difference becomes much smaller. We believe that the variation in log-norm distributions may be caused by the model’s varying degrees of fit to different categories. Although there is significant overlap in the norms across different categories and sessions, preventing direct prediction based solely on norms, norms can still serve as a part of the basis for classification.
Figure 4: Histogram of the distribution of the log norms of sample embedding on miniImageNet dataset. (a) The base session. (b) 2 incremental sessions. (c) 3 randomly selected classes in the base session. (d) 3 randomly selected classes in incremental sessions. The curve represents the normal distribution formed by the sample mean and variance of the log norms.
@@ 4.2.2 Dual-Granularity parametric estimation
As shown in Fig. 4, we observe significant differences between the incremental and base sessions, but the distributions across different incremental sessions are approximately similar. Based on this phenomenon, we choose to estimate a distribution for the norms of each class. Based on the trend of class distributions in Fig. 4(b) and 4(d), we assume that the log norms of the embeddings for samples belonging to the same category follow a normal distribution, represented as $\ln{\|\mathbf{e}_{i}\|}\sim\mathcal{N}(\mu_{j},\sigma^{2}_{j})$, where $y_{i}=j$, $j$ is the category label, $\mu_{j}$ denotes the mean, and $\sigma^{2}_{j}$ denotes the variance.
During the base session, with sufficient training samples, we approximate the population mean and variance with sample mean and variance, respectively. As shown in Fig. 4(b), the differences in log norms between classes are significant, making it necessary to perform class-level modeling of the log norms. During the incremental sessions, however, each class contains only a few samples, which may result in significant deviations between the sample mean and variance and the true population values. Despite this, as shown in Fig. 4(d), the differences in log norm between classes during these sessions are actually quite small. We assume that the log norm of each incremental category follows the same normal distribution. We know that if the distribution of samples in each dataset follows the same normal distribution, then mixing the datasets will result in the samples still following a normal distribution, denoted as $\ln{\|\mathbf{e}_{i}\|}\sim\mathcal{N}(\mu_{\text{incre}},\sigma^{2}_{\text{incre}})$, where $y_{i}\in(\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t})$, $t\geq 1$, and $\mu_{\text{incre}},\sigma^{2}_{\text{incre}}$ are the shared mean and variance, respectively. We conduct a test in Sec. 5.5 to determine whether the classes follow a normal distribution. We refer to assigning the same distribution to different classes in the incremental sessions as the shared distribution technique. A session-granular distribution does not directly improve classification accuracy. Its primary purpose is to distinguish between the base session and incremental sessions. By improving the accuracy of session identification, it indirectly enhances the final classification accuracy. In summary, we model the distributions of the base and incremental rounds with different levels of granularity.
@@ 4.2.3 Angle-Norm joint logits inference
The pipeline of Angle-Norm joint (ANJ) inference is illustrated in Fig. 3. ANJ consists of 2-Stage NCM (2SNCM) and Norm Distribution.
@@
where $\hat{\mathbf{e}}_{i}=\frac{\mathbf{e}_{i}}{\|\mathbf{e}_{i}\|}$. During the base session, we normalize embeddings to treat each sample equally, because otherwise, samples with larger norms would have greater weight in the calculation of $\boldsymbol{w}_{j}$. As illustrated in Fig. 4(a), there is considerable variance in the norms of samples during the base session. In the incremental sessions, due to the scarcity of samples, normalizing embeddings could lead to a significant loss of information. Therefore, we adopt a staged approach to better accommodate FSCIL.
@@
where $X$ represents a random variable that follows the distribution of norms for class $j$. We can obtain the probability by the cumulative distribution function of normal distribution. We also explore other distributions in Sec. 5.4.
Angle-Norm Joint Inference. Since norms contain less information than angles, we first compress the norm logits, ultimately obtaining the joint logits. The final angle-norm joint logits $\mathbf{z}_{j}$ is calculated by $\mathbf{z}_{j}=\mathbf{z}_{j}^{1}\left(\mathbf{z}_{j}^{2}\right)^{C}$, where $C$ is the compression coefficient. After compression, the differences between the norm logits of different categories are reduced, thereby amplifying the role of the angle logits. The final predicted category is determined by $\mathbf{z}_{j}$.
@@ 5.1 Experimental Setup
Datasets. Following the previous works setting [8, 11, 9, 16], we perform conventional experiments on three datasets: CIFAR100 [50], CUB200 [56] and miniImageNet [57]. To better reflect real-world situations, we also conduct open-ended and imbalanced experiments on the CompCars [58] dataset. CIFAR100 consists of 100 different classes, with each class containing 600 images. CUB200 contains images of 200 bird species. miniImageNet is a subset of ImageNet [1] and consists of 100 selected classes. CompCars is a fine-grained vehicle dataset that contains 431 categories, and the large number of categories allows us to set up long-term learning experiments to simulate open-ended scenarios. In CIFAR100 and miniImageNet, the base session consists of 60 classes, while the incremental sessions comprise 40 classes. During the incremental sessions, the data is organized in a 5-way 5-shot format. As for CUB200, the base session includes 100 classes, and the incremental sessions involve 100 classes as well. In the incremental sessions of CUB200, the data is organized in a 10-way 5-shot format.
Implementation Details. We conduct experiments by separately employing SAAN, and embed it into SOTA methods, $i.e.$ FACT [11], SAVC [16]. Both of these methods adopt the feature freezing paradigm and use NCM classifier. For fairness, all these methods adopt ResNet18 [2] for miniImageNet, CUB200 and CompCars, and ResNet20 [2] for CIFAR100. In SAAN, the initial learning rate is set to 0.1 with cosine annealing on CIFAR100,miniImageNet, and milestone decay on CUB200 and CompCars. The batch size is 256, and training is conducted for 600 epochs. The training parameters in the embedding of FACT and SAVC are set to be the same as those in their works. The selection of hyper-parameters used in SAAN is discussed in Sec. 5.5.
Table 1: Performance comparison on CUB200 dataset. *: FSCIL results implemented by [8]. $\ddagger$:reproduced under the FSCIL setting with the source code. Drop: The drop in Acc. of the last session relative to session 0.
@@
Table 2: Performance comparison on miniImageNet dataset. *: FSCIL results implemented by [8]. $\ddagger$:reproduced under the FSCIL setting with the source code. Drop: The drop in Acc. of the last session relative to session 0.
@@
Table 3: Performance comparison on CIFAR100 dataset. *: FSCIL results implemented by [8]. $\ddagger$:reproduced under the FSCIL setting with the source code. Drop: The drop in Acc. of the last session relative to session 0.
@@ 5.2 Performance on Conventional Benchmarks
Compared to Decoupled-Cosine (frozen paradigm baseline), our method exhibits an average improvement over 7.44% across the three datasets in the final session, highlighting the importance of prior spatial allocation. It can be observed that SAAN surpasses FACT on all three datasets, indicating that our method achieves the performance of SOTA models. FACT occupies the sample space with virtual classes, but new classes may not appear within the space occupied by these virtual classes. SAAN further aligns the virtual classes semantically with the new classes, addressing the limitations of FACT, which allows SSAN to achieve better performance.
SAAN does not outperform SAVC, primarily because SAVC not only uses the virtual class method but also incorporates supervised contrastive learning. Based on the self-supervised MoCo [61] framework, it significantly increases both data volume and training capacity with additional supervisory signals, yielding highly effective results. The generalization capability provided by contrastive learning is crucial for more complex datasets. However, on simpler datasets like CIFAR100, SAAN already outperforms SAVC. This is because SAAN requires generalized features. In simple datasets, the model can easily obtain relatively generalized features, but in more challenging datasets, additional assistance, such as supervised contrastive learning, is needed.
@@
SAAN shows significant performance advantages on CUB200 and achieves good results on miniImageNet. However, it performs worse than NC-FSCIL on CIFAR100. One possible explanation is that in SAAN, prototype allocation considers semantic information of the samples, and prototypes are adjustable. This gives SAAN good performance on fine-grained datasets like CUB200, where certain bird species have high similarity, and the similarity between different classes is more important to consider. In contrast, on datasets like CIFAR100, where there is a larger difference between categories, the differences between each category are relatively large, so the importance of carefully adjusting and allocating prototypes may be reduced. Instead, the random allocation and fixed strategy in NC-FSCIL ensure the stability of prototypes during the incremental process, leading to the best performance.
Table 4: The experimental results under open-ended and imbalanced conditions on CompCars dataset. LA: the Last session Accuracy(%), AA: the Average Accuracy (%), $\Delta_{last}$: Relative improvements of the last sessions compared to the Decoupled-Cosine, $\Delta_{average}$: Relative improvements of the last sessions compared to the Decoupled-Cosine, Drop: The drop in accuracy of the last session relative to session 0.
Method LA $\uparrow$ $\Delta_{last}$ AA $\uparrow$ $\Delta_{average}$ Drop $\downarrow$
Decoupled-Cosine[60] 56.24 69.23 29.76
SAAN 62.27 +6.03 76.24 +7.01 31.38
FACT[11] 61.55 +5.31 75.23 +6.00 31.23
FACT+SAAN 63.59 +7.35 76.52 +7.29 29.60
SAVC [16] 70.76 +14.52 81.32 +12.09 24.24
SAVC+SAAN 73.59 +17.35 83.33 +14.10 21.93
5.3 Comparisons in Open-Ended and Imbalanced Scenarios
To better simulate real-world class-incremental task scenarios, we conduct the experiments under open-ended and imbalanced conditions. Unlike the experiments in Sec. 5.2, we use the CompCars dataset, a fine-grained vehicle dataset with 431 car categories. Among them, 231 classes are used for the base session, while 200 classes are allocated for the incremental sessions. To simulate long-term few-shot class-incremental tasks, we extend the number of sessions to 21, including 20 incremental sessions. Additionally, the number of categories in each incremental session is sampled from a Gaussian distribution with a mean of 10 and a variance of 5, rather than being fixed at 10 classes. Similarly, the number of samples per category is sampled from a Gaussian distribution with a mean of 5 and a variance of 2, rather than being fixed at 5 samples. This setup is designed to simulate the class and sample imbalance characteristics in the real world. Due to the randomness in the task setup, we conduct 3 repeated experiments based on this configuration, and all experimental data are presented as the average of repeated experiments.
We perform experiments on Decoupled-Cosine, FACT, SAVC and SAAN, and integrate our method as a plug-in into FACT and SAVC, like Sec. 5.2. To ensure fairness, all methods are implemented using the same random seeds. We show the experiment results in Tab. 4, and the detailed results are provided in C. Here, the Last session Accuracy (LA), the Average Accuracy of all sessions (AA) and The drop in accuracy of the last session relative to session 0 (Drop) are used as evaluation metrics.
The experimental results are largely consistent with those in Sec. 5.2, particularly resembling the outcomes on CUB200 in Tab. 1. This aligns with our expectations, as both being fine-grained datasets, CompCars poses similar challenges to these methods as CUB200 does. SAAN shows an improvement of 6.03% in LA and 7.01% in AA over decoupled-cosine (the frozen paradigm baseline), and an enhancement of 0.72% in LA and 1.01% in AA over FACT, indicating that our method can achieve the performance of SOTA models when used independently, and remains effective in open-ended and imbalanced scenarios. SAAN does not outperform SAVC, which is also in line with our previous analysis. Compared to coarse-grained datasets with distinct inter-class differences such as CIFAR100, models require greater generalization capability on fine-grained datasets, and the supervised contrastive learning used in SAVC can well compensate for this. Therefore, when we integrate SAAN as a plugin into SAVC, the model’s performance significantly surpasses that of either SAAN or SAVC alone, with a 2.83% improvement in LA and a 2.01% improvement in AA over SAVC, achieving the highest performance and the lowest forgetting rate. SAAN enhances the performance of SAVC through two key refinements: it meticulously adjusts the spatial arrangement of SAVC’s semantic virtual categories and enhances the NCM classifier within SAVC to more effectively utilize the angular and magnitude data derived from a constrained set of samples. Embedding SAAN into FACT also boosts FACT’s performance. This demonstrates that SAAN remains effective as a plugin in open-ended and imbalanced scenarios.
@@ 5.3 Comparisons in Open-Ended and Imbalanced Scenarios
Table 6: ANJ ablation study on CUB200 dataset.
Distribution Shared Distribution Acc. in each session (%$\uparrow$
0 1 2 3 4 5 6 7 8 9 10
Normal $\times$ 78.19 73.84 70.29 66.36 65.68 62.11 61.55 60.10 59.69 58.20 57.56
$\checkmark$ 78.19 74.56 71.14 66.79 66.29 63.02 62.03 60.87 60.01 59.15 58.18
Log-Normal $\times$ 78.42 74.03 70.63 66.39 65.90 62.52 61.82 60.44 59.66 58.72 57.72
$\checkmark$ 78.42 74.94 71.54 67.17 66.67 63.38 62.43 61.27 60.49 59.55 58.55
Weibull $\times$ 78.43 74.48 70.90 66.75 65.94 62.80 61.63 60.18 59.68 58.96 57.70
$\checkmark$ 78.43 74.98 71.58 67.30 66.62 63.48 62.47 61.16 60.51 59.71 58.65
Patero $\times$ 76.44 72.03 68.33 64.30 63.85 60.09 60.02 58.62 57.61 56.86 55.46
$\checkmark$ 76.44 72.30 68.62 64.70 64.14 60.45 60.48 58.95 57.99 57.19 55.76
@@ 5.4 Ablation Studies
The experiments reveal that $\mathcal{L}_{1}$ has the most significant impact, with a 3.41% improvement when combined with the baseline, highlighting the importance of embedding space allocation. $\mathcal{L}_{2}$ aims to increase inter-class distances, resulting in a 4.01% improvement. 2SNCM achieves a modest improvement of 4.19%, indicating a limited enhancement. This indicates that 2SNCM can achieve more reasonable representative points simply by normalizing and averaging the samples in the base round, ultimately leading to a certain performance improvement. To further verify the effectiveness of 2SNCM, we present multiple repeated experiments on three datasets in E. Norm Distribution achieves an improvement of 5.18%, demonstrating the effectiveness of introducing norm information to the classifier. The ANJ, composed of 2SNCM and Norm Distribution, ultimately achieves a performance improvement of 5.44%. The components in ANJ do not participate in model training but construct a more accurate classifier based on the feature extractor, with minimal computational cost. In summary, the spatial allocation strategy of CCSA is highly beneficial for FSCIL, while ANJ aids classification by making more comprehensive and effective use of the features.
To verify the rationality of the distribution assumption in ANJ and the effectiveness of the shared distribution technique, we conduct an ablation study on ANJ alone, as shown in Tab. 6. The log-normal, normal, and Weibull distributions can effectively represent feature norms, leading to a certain improvement in model performance. However, the Pareto distribution, a heavy-tailed distribution, rendered ANJ ineffective due to its significant difference from the norm distribution. Among all distributions, the Weibull and log-normal distributions performed the best. The log-normal distribution has the advantage of simple parameter estimation, whereas the Weibull distribution requires storing data for distribution fitting. Without adopting the shared distribution technique, the incremental data becomes too sparse, leading to a significant discrepancy between the estimated distribution and the true distribution. This issue is particularly severe for the Weibull distribution, which requires a sufficient amount of data for proper fitting. Regardless of the chosen distribution, without the shared distribution technique, the performance deteriorates significantly. By leveraging the shared distribution technique, the model can better identify the session to which a sample belongs, thereby indirectly improving the final classification performance.
@@
Figure 5: The t-SNE visualization on CIFAR100 dataset of the embeddings. Classes 0-5 represent the base classes while classes 6-8 represent the novel classes. The squares represent the centers set by SAAN.